OpenStudy (batmaaaanxd):

Find the zeroes of the polynomial 2x-5 and verify the relationship between zeroes and coefficients.

OpenStudy (batmaaaanxd):

Please help, idk how to solve without splitting the middle term and it cannot be applied here.

OpenStudy (batmaaaanxd):

@phi

OpenStudy (phi):

"find the zeros" means find where 2x-5= 0 you should "solve for x" start by add +5 to both sides

OpenStudy (batmaaaanxd):

so x = 5/2? @phi

OpenStudy (batmaaaanxd):

the zeroes will be -5/2 and 5/2, right? And to verify we have to do that alpha + beta = -b/a and alpha * beta = c/a, right? @phi